Hi. My teacher gave me a packet to do for a test next wed and i was wondering if you could help me with a problem.
Solve give completely simplified answers in terms of natural logs.
3^x = 5^x-1
thanks for your help

Answer by Alan3354(69443)

Solve give completely simplified answers in terms of natural logs.
3^x = 5^x-1
--------------
Assuming x-1 is the exponent of 5:
x*log(3) = (x-1)*log(5)
x*log(3) = x*log(5) - log(5)
x(log(3) - log(5)) = -log(5)
x = log(5)/(log(5) - log(3))
just calculator work from there
-----
Ooh, natural logs
x = ln(5)/(ln(5) - ln(3))